YBR:

What value of b will cause the system to have an infinite number of solutions? y = 6x – b -3x+ 1/2y= -3 b = a 2 b 4 c 6 d 8

jhonyy9:

y = 6x -b -3x +(1/2)y = -3 do you know a system when has an infinite number of solutions ?

YBR:

huh?

jhonyy9:

simplifie the second equation in the form of the first equation

jhonyy9:

i like help you please collaborate

jhonyy9:

so like a first step add to both sides 3x what will get in this way ?

YBR:

x+1/2y=-3?

jhonyy9:

no then you add to -3x +3x what will get ?

jhonyy9:

and i ve said add to both sides 3x

YBR:

0

jhonyy9:

yes exactly so please rewrite this all

jhonyy9:

-3x +3x +(1/2)y = -3 +3x

jhonyy9:

continue please

YBR:

1/2y

jhonyy9:

yes on the left side and on the right side ? please write the equation

YBR:

bro nvm cuh I gotta quiz and you sitting up here tryna do extra

jhonyy9:

ok but i not can give you never direct answer - i like teach you how you can solve the same math problems have a nice day bye bye

YBR:

mhm, mkay